# Nonstandard Lagrangians and branched Hamiltonians: A brief review

### v1

@@ Abstract
Time and again, non-conventional forms of Lagrangians have found attention in the literature. For one thing, such Lagrangians have deep connections with several aspects of nonlinear dynamics including specifically the types of the Liénard class; for another, very often the problem of their quantization opens up multiple branches of the corresponding Hamiltonians, ending up with the presence of singularities in the associated eigenfunctions. In this article, we furnish a brief review of the classical theory of such Lagrangians and the associated branched Hamiltonians, starting with the example of Liénard-type systems. We then take up other cases where the Lagrangians depend upon the velocity with powers greater than two while still having a tractable mathematical structure, while also describing the associated branched Hamiltonians for such systems. For various examples, we emphasize upon the emergence of the notion of momentum-dependent mass in the theory of branched Hamiltonians.
@@ I Introduction
During the past few years, the study of non-conventional types of dynamical systems, in particular those which are controlled by Lagrangians that are not quadratic in the velocity has entered a new phase of intense development SW ; SW1 ; SW2 ; MCZ . Such Lagrangians lead to certain exotic Hamiltonians, commonly termed as branched Hamiltonians, that have relevance in their applicability to problems of nonlinear dynamics pertaining to autonomous differential equations bspf ; mitso , and to certain exotic quantum-mechanical models especially in the context of non-hermitian parity-time ($\mathcal{PT}$)-symmetric schemes bender , along with their relativistic counterparts bpm .
A simple way to see how Lagrangians that are not quadratic in the velocity can lead to sensible dynamical systems is to consider the following toy model benoy ; carinena :
@@
where $\alpha$ and $\beta$ are real numbers satisfying $\alpha\beta>0$, and $\alpha x+\beta\dot{x}\neq 0$. Notice that the Lagrangian cannot be expressed as the difference between the kinetic and potential energies; such Lagrangians shall be referred to as nonstandard. A direct computation reveals that the Euler-Lagrange equation is
@@
where $\gamma=\frac{3\alpha}{2\beta}$ and $\omega_{0}=\frac{\alpha}{\sqrt{2}\beta}$. Eq. (2) is just the harmonic oscillator in the presence of linear damping. We remind the reader that there is no time-independent Lagrangian of the ‘standard’ kind from which one can reproduce Eq. (2) upon invoking the Euler-Lagrange equation [^4] One could recover the damped oscillator from a standard Lagrangian by using a Rayleigh dissipation function goldsteincm . Alternatively, one can consider the modified Euler-Lagrange equations from the Herglotz variational problem to describe the damped oscillator herglotz . We do not consider such situations here.. There exist various other families of nonstandard Lagrangians (giving rise to different dynamical systems) which look quite different from Eq. (1); each family is endowed with their own intriguing features. However, the common theme is the existence of Lagrangians that are not quadratic in the velocity, thereby leading to a nonlinear relationship between the velocity and the momentum. This leads to the notion of branching, typically yielding the so-called Riemann-surface phase-space structure and in consequence, certain interesting topological features CZ1 .
@@
Against this background, a new class of innovations on the description and simulations of quantum dynamics emerged in relation to the specific role played by certain models constructed appropriately. Not quite unexpectedly, Hamiltonians which are multi-valued functions of momenta confront us with some typical insurmountable ambiguities of quantization. In such cases, the underlying Lagrangian possesses time derivatives in excess of quadratic powers (or sometimes, inverse powers). The use of these models leads, on both classical and quantum grounds, to the necessity of a re-evaluation of the dynamical interpretation of the momentum, which, in principle, becomes a multi-valued function of velocity. It also needs to be pointed out that the traditional approaches often do not always work as is the case with certain $\mathcal{PT}$-symmetric complex potentials possessing real spectra BagG or, on employing tractable non-local generalizations zno1 .
@@
There have been efforts to construct simple systems with toy Lagrangians that lead to multi-valued Hamiltonian systems. Such Hamiltonians would be defined in the momentum space revealing systems with momentum-dependent masses BGG ; agp ; brp . In the context of nonlinear models, the Liénard system presents an intriguing feature of the Hamiltonian in which the roles of the position and momentum variables often get exchanged BGG ; ML . Thus, Liénard systems are of potential importance in optics DOR as well as in non-Hermitian quantum mechanics BGG ; Bag2 . Naturally, the presence of the damping as is the case for Liénard systems poses to be a problem whenever one tries to contemplate a quantization of the model. It is important to realize that the quantization is hard to tackle in the coordinate representation of the Schrödinger equation but can be straightforwardly carried out in the momentum space. Such a Hamiltonian, which has a branched character, may be interpreted to represent an effective-mass quantum system BGG ; Roos ; ruby .
@@
Although much has been said about the quantum-mechanical formalisms, in this paper, we focus on the classical theory, (briefly) reviewing some aspects of nonstandard Lagrangians and the associated branched Hamiltonians. The theory is exemplified by focusing on various examples which include some systems of the Liénard class, which are of great interest in the theory of dynamical systems. Apart from Liénard systems, we discuss some interesting toy Lagrangians which contain time derivatives in excess of quadratic powers, leading to branched Hamiltonians. The basic features of the theory are discussed in the light of these examples. However, we begin with a discussion on some simple nonstandard Lagrangians which can be figured out via some guesswork; we call them trial Lagrangians and they are discussed in Section II. Following this, in Section III we discuss nonstandard Lagrangians and branched Hamiltonians in the context of Liénard systems, wherein we outline a systematic derivation of the Lagrangians, provided the system admits a certain integrability condition. In Sections IV, V and VI we analyze various intriguing examples of Lagrangians in which time derivatives occur in excess of quadratic powers, while also discussing the associated Hamiltonians. We conclude with some remarks in Section VII where, in particular, a few further aspects of the problem of quantization are discussed. .
II Trial Lagrangians: Reciprocal and logarithmic kinds
Sometimes one can make educated guesses to construct plausible nonstandard Lagrangians. We cite a few examples.
@@ II Trial Lagrangians: Reciprocal and logarithmic kinds
where $\mu(x)$ is a well-behaved function (typically a polynomial), while $\alpha$ and $\beta$ are real-valued and non-zero constant numbers. Obviously, it does not reveal the ‘standard’ form as the difference between the kinetic and potential energies. However, the Euler-Lagrange equation gives $\ddot{x}+f(x)\dot{x}+g(x)=0$, with $f(x)=\frac{3\alpha\mu^{\prime}(x)}{2\beta}$ and $g(x)=\frac{\alpha^{2}\mu^{\prime}(x)\mu(x)}{2\beta^{2}}$, where for instance, picking $\mu(x)=x$ gives the linearly-damped harmonic oscillator, while the choice $\mu(x)=x^{2}$ implies $f(x)\propto x$ and $g(x)\propto x^{3}$. Lagrangians of this type [Eq. (3)] are termed as reciprocal Lagrangians.
@@
$L(x,\dot{x})=\ln[\gamma\mu(x)+\delta\dot{x}],\quad\quad\gamma\mu(x)+\delta\dot{x}\neq 0,$ (4)
where $\delta$ and $\beta$ are real-valued and non-zero constant numbers. The Euler-Lagrange equation goes as $\ddot{x}+f(x)\dot{x}+g(x)=0$, with $f(x)=\frac{2\gamma\mu^{\prime}(x)}{\delta}$ and $g(x)=\frac{\gamma^{2}\mu(x)\mu^{\prime}(x)}{\delta^{2}}$. Lagrangians given by Eq. (4) are termed as logarithmic Lagrangians.
@@
However, there is a limited variety of differential equations that can be described by Lagrangians which may be guessed; in general, it is often not possible to systematically derive a Lagrangian from which the equation may emerge as the Euler-Lagrange equation. In what follows, we describe Liénard systems and demonstrate that if a certain integrability condition is satisfied, then one may systematically find nonstandard Lagrangians describing such systems.
@@ A Liénard system is a second-order ordinary differential equation that goes as
where [^5] Often it is sufficient to have $f(x),g(x)\in C^{2}(\mathbb{R},\mathbb{R})$. $f(x),g(x)\in C^{\infty}(\mathbb{R},\mathbb{R})$ can be suitably chosen. Interesting choices for $f(x)$ and $g(x)$ include $f(x)=1$ and $g(x)=x$, which is just the damped linear oscillator, while the choice $f(x)=(1-x^{2})$ and $g(x)=x$ gives the van der Pol oscillator, known to admit limit-cycle behavior due to the particular choice of $f(x)$ str . Another choice is $f(x)=1$ and $g(x)=x^{3}$, for which we have the linearly-damped (nonlinear) Duffing oscillator. It is noteworthy that in any case with $f(x)\neq 0$, the system exhibits non-conservative dynamics because Eq. (8) does not stay invariant under the transformation $t\rightarrow-t$, namely, time reversal. Further, oscillatory dynamics can be obtained if $f(x)$ is an even function and if $g(x)$ is odd; it follows from the fact that the overall force (the second and third terms of Eq. (8)) should be odd under $x\rightarrow-x$ in order to support oscillations mic .
@@ III.1 Cheillini condition and nonstandard Lagrangians
We now discuss a systematic derivation of nonstandard Lagrangians describing certain Liénard systems agp ; BGG . The idea is to make use of a function known as the Jacobi last multiplier which may be defined as follows whittaker . Given a second-order ordinary differential equation,
@@
one defines the last multiplier $M$ as a solution of the following differential equation:
@@
As has been discussed in Whittaker’s classic textbook whittaker , if a second-order differential equation such as Eq. (9) follows from the Euler-Lagrange equations, then the Lagrangian is related to the last multiplier as
$M=\frac{\partial^{2}L}{\partial\dot{x}^{2}}.$ (11)
This allows one to determine the Lagrangian function for a given second-order differential equation, provided it admits a Lagrangian formalism. For the Liénard system, a formal solution for the last multiplier is found to be
@@
where $u$ is some nonlocal variable and $\ell$ is a suitable parameter whose significance appears in the following lemma:
Lemma 1
The Liénard equation [Eq. (8)] can be written as the following system of first-order equations:
$\dot{u}=\ell uf(x),\;\;\;\dot{x}=u+W(x),$ (13)
where $W(x)=\ell^{-1}g(x)/f(x)$, with the parameter $\ell$ being determined by the following condition:
@@
Proof - From Eq. (12) we have
[…]
which means $\dot{u}=\ell uf(x)$. Simply putting $\dot{x}=u+W(x)$, we find by differentiating with respect to $t$, the following expression:
$\ddot{x}=\dot{u}+W^{\prime}(x)\dot{x}.$ (16)
Thereafter, substituting the expression for $\dot{u}$ from the previous equation and after eliminating $u$, it follows that
$\ddot{x}=\ell f(x)(\dot{x}-W(x))+W^{\prime}(x)\dot{x}.$ (17)
Comparison with Eq. (8) reveals that one must choose $W^{\prime}(x)=-(\ell+1)f(x)$ and $W(x)=\ell^{-1}g(x)/f(x)$. Consistency between these two choices therefore demands that the condition appearing in Eq. (14) should be fulfilled.
It is noteworthy that Eq. (14) represents what is known as the Cheillini condition, allowing one to recast the Liénard system in the form suggested in Eq. (13). Specifically, if $f(x)=ax^{\alpha}$, then Eq. (14) dictates that $g(x)$ must satisfy the following differential equation:
@@
where $k\in\mathbb{R}$ is an integration constant. This gives the functional form of $g(x)$ so as to satisfy the Cheillini condition. In the cases where the Cheillini condition is satisfied, one can find the Lagrangian as follows:
Theorem 1
Consider a Liénard system with choices of $f(x)$ and $g(x)$ that satisfy the Cheillini condition. Then, the second-order differential equation can be obtained as the Euler-Lagrange equation of the Lagrangian that reads
$L(x,\dot{x})=\frac{\ell^{2}}{(\ell+1)(2\ell+1)}\left(\dot{x}-\frac{1}{\ell}\frac{g(x)}{f(x)}\right)^{{(2\ell+1)}/{\ell}}.$ (20)
Proof - If the functions $f(x)$ and $g(x)$ satisfy Eq. (14), then we have from Eqs. (11) and (12), the following condition:
$\frac{\partial^{2}L}{\partial\dot{x}^{2}}=\left(\dot{x}-\frac{1}{\ell}\frac{g(x)}{f(x)}\right)^{1/\ell}.$ (21)
Then, it follows straightforwardly by simple integration that the Lagrangian is given by Eq. (20). Clearly, a Lagrangian such as the one appearing in Eq. (20) is of the nonstandard form because there is no identification of kinetic and potential energies. It may be emphasized that not all choices of $f(x)$ and $g(x)$ would satisfy the Cheillini condition, meaning that such equations cannot be cast into the form of Eq. (13); an example is the van der Pol oscillator for which $f(x)=(1-x^{2})$ and $g(x)=x$. Therefore, not all Liénard systems are known to admit a Lagrangian description.

### v2

# A reappraisal of Lagrangians with non-quadratic velocity dependence and branched Hamiltonians

@@ Abstract
Time and again, non-conventional forms of Lagrangians with non-quadratic velocity dependence have found attention in the literature. For one thing, such Lagrangians have deep connections with several aspects of nonlinear dynamics including specifically the types of the Liénard class; for another, very often the problem of their quantization opens up multiple branches of the corresponding Hamiltonians, ending up with the presence of singularities in the associated eigenfunctions. In this article, we furnish a brief review of the classical theory of such Lagrangians and the associated branched Hamiltonians, starting with the example of Liénard-type systems. We then take up other cases where the Lagrangians depend upon the velocity with powers greater than two while still having a tractable mathematical structure, while also describing the associated branched Hamiltonians for such systems. For various examples, we emphasize upon the emergence of the notion of momentum-dependent mass in the theory of branched Hamiltonians.
@@ I Introduction
During the past few decades, the study of non-conventional types of dynamical systems, in particular those which are controlled by Lagrangians that are not quadratic in the velocity has entered a new phase of intense development SW ; SW1 ; SW2 ; MCZ . Such Lagrangians lead to certain exotic Hamiltonians, commonly termed as branched Hamiltonians, that have relevance in their applicability to problems of nonlinear dynamics pertaining to autonomous differential equations bspf ; mitso , and to certain exotic quantum-mechanical models especially in the context of non-hermitian parity-time ($\mathcal{PT}$)-symmetric schemes bender , along with their relativistic counterparts bpm .
A simple way to see how Lagrangians that are not quadratic in the velocity can lead to meaningful dynamical systems is to consider the following toy model benoy ; carinena (see also, Refs. CFN ; CGR09 ):
@@
where $\alpha$ and $\beta$ are real numbers satisfying $\alpha\beta>0$. We may also require that $\alpha x+\beta\dot{x}\neq 0$, i.e., that the velocity phase space accessible to the system is defined as a subset of $\mathbb{R}^{2}$.
Notice that the Lagrangian cannot be expressed as the difference between the kinetic and potential energies; such Lagrangians shall be referred to as nonstandard, i.e., in this paper we will be adopting such a nomenclature in which the term ‘nonstandard Lagrangian’ would refer to a Lagrangian with a non-quadratic velocity dependence [^5] A linear dependence on the velocity makes the Hessian matrix singular resulting in a singular Legendre transform while passing from the Lagrangian to the Hamiltonian formalism (see for example, Ref. dirac ). We do not address such cases here and deal with Lagrangians that have velocity dependence either in excess of quadratic powers or in inverse powers..
A direct computation reveals that the Euler-Lagrange equation is
@@
where $\gamma=\frac{3\alpha}{2\beta}$ and $\omega_{0}=\frac{\alpha}{\sqrt{2}\beta}$. (2) is just the harmonic oscillator in the presence of linear damping. We remind the reader that there is no time-independent Lagrangian of the ‘standard’ kind from which one can reproduce (2) upon invoking the Euler-Lagrange equation [^6] One could recover the damped oscillator from a standard Lagrangian by using a Rayleigh dissipation function goldsteincm . Alternatively, one can consider the modified Euler-Lagrange equations from the Herglotz variational problem to describe the damped oscillator herglotz . We do not consider such situations here.. There exist various other families of nonstandard Lagrangians (giving rise to different dynamical systems) which look quite different from (1); each family is endowed with their own intriguing features. However, the common theme is the existence of Lagrangians that are not quadratic in the velocity, thereby leading to a nonlinear relationship between the velocity and the momentum. It may be emphasized that a Lagrangian $L=L(x,\dot{x})$ defined for a system whose configuration space is a subset of $\mathbb{R}$ is called regular if its Hessian with respect to the velocity is non-vanishing, i.e.,
$\frac{\partial^{2}L(x,\dot{x})}{\partial\dot{x}^{2}}\neq 0,$ (3)
and is of constant sign allowing us to solve for the velocity $\dot{x}$ in favor of the momentum $p(x,\dot{x})=\frac{\partial L(x,\dot{x})}{\partial\dot{x}}$, i.e., we can write $\dot{x}(x,p)$. Thus, Lagrangians with a quadratic velocity dependence are regular and one can formulate a Hamiltonian description by the means of a Legendre transform. The condition (3) fails for Lagrangians that are linear in the velocity (see for instance, Ref. dirac ) but as mentioned earlier they won’t be our concern here. Instead, we shall be looking at Lagrangians for which solving the equation $p=(x,\dot{x})$ in favor of the $\dot{x}$ leads to a non-unique solution, e.g., the appearance of a square root which gives rise to what will be called branching. Such Lagrangians would not permit the construction of a Hamiltonian function in a unique way.
@@
Against this background, a new class of innovations on the description and simulations of quantum dynamics emerged in relation to the specific role played by certain models constructed appropriately. Not quite unexpectedly, Hamiltonians which are multi-valued functions of momenta confront us with some typical insurmountable ambiguities of quantization. In such cases, the underlying Lagrangian possesses time derivatives in excess of quadratic powers (or sometimes, inverse powers). The use of these models leads, on both classical and quantum grounds, to the necessity of a re-evaluation of the dynamical interpretation of the momentum, which, in principle, becomes a multi-valued function of the velocity. It also needs to be pointed out that the traditional approaches often do not always work as is the case with certain $\mathcal{PT}$-symmetric complex potentials possessing real spectra BagG or, on employing tractable non-local generalizations zno1 .
@@
In the context of nonlinear models, certain Liénard-class systems present an intriguing feature of the Hamiltonian in which the roles of the position and momentum variables get exchanged with the emergence of the notion of a momentum-dependent mass agp ; BGG ; brp ; ML ; Bag2 ; ruby ; modak . Naturally, the presence of the damping as is the case for Liénard systems poses to be a problem whenever one tries to contemplate a quantization of the model. It is important to realize that the quantization is hard to tackle in the coordinate representation of the Schrödinger equation but can be straightforwardly carried out in the momentum space BGG ; ruby (see also, Ref. Roos ).
@@
Although much has been said about the quantum-mechanical formalisms, in this paper, we focus on the classical theory, (briefly) reviewing some aspects of nonstandard Lagrangians and the associated branched Hamiltonians. The theory is exemplified by focusing on various examples which include some systems of the Liénard class, which are of great interest in the theory of dynamical systems. Apart from Liénard systems, we discuss some interesting toy Lagrangians which contain time derivatives in excess of quadratic powers, leading to branched Hamiltonians. The basic features of the theory are discussed in the light of these examples. However, we begin with a discussion on some simple nonstandard Lagrangians which can be figured out via some guesswork in Section II. Following this, in Section III we discuss nonstandard Lagrangians and branched Hamiltonians in the context of Liénard systems, wherein we outline a systematic derivation of the Lagrangians, provided the system admits a certain integrability condition. In Sections IV and V we analyze various intriguing examples of Lagrangians in which time derivatives occur in excess of quadratic powers, while also discussing the associated Hamiltonians. We conclude with some remarks in Section VI and in Appendix A where a few further aspects of the problem of quantization are also discussed.
II Some illustrative examples
@@
where $\mu(x)$ is a well-behaved function (typically a polynomial), while $\alpha$ and $\beta$ are real-valued and non-zero constant numbers. Obviously, it does not reveal the ‘standard’ form as the difference between the kinetic and potential energies. However, the Euler-Lagrange equation gives $\ddot{x}+f(x)\dot{x}+g(x)=0$, with $f(x)=\frac{3\alpha\mu^{\prime}(x)}{2\beta}$ and $g(x)=\frac{\alpha^{2}\mu^{\prime}(x)\mu(x)}{2\beta^{2}}$, where for instance, picking $\mu(x)=x$ gives the linearly-damped harmonic oscillator, while the choice $\mu(x)=x^{2}$ implies $f(x)\propto x$ and $g(x)\propto x^{3}$. Lagrangians of this type [(4)] are termed as reciprocal Lagrangians.
@@
$L(x,\dot{x})=\ln[\gamma\mu(x)+\delta\dot{x}],\quad\quad\gamma\mu(x)+\delta\dot{x}>0,$ (5)
where $\delta$ and $\beta$ are real-valued and non-zero constant numbers. The Euler-Lagrange equation goes as $\ddot{x}+f(x)\dot{x}+g(x)=0$, with $f(x)=\frac{2\gamma\mu^{\prime}(x)}{\delta}$ and $g(x)=\frac{\gamma^{2}\mu(x)\mu^{\prime}(x)}{\delta^{2}}$. Lagrangians that look like (5) are termed as logarithmic Lagrangians. The relation between logarithmic and reciprocal classes of Lagrangians has been explored in CFN (see also, Ref. CGR09 ) [^7] As with the system described by the Lagrangian (1), the systems given by (4) and (5) are defined only on appropriate regions of $\mathbb{R}^{2}$..
@@
However, there is a limited variety of differential equations that can be described by Lagrangians which may be guessed; in general, it is often not possible to systematically derive a Lagrangian from which a given differential equation may emerge as the Euler-Lagrange equation. In what follows, we describe Liénard systems and demonstrate that if a certain integrability condition is satisfied, then one may systematically find nonstandard Lagrangians describing such systems.
@@ A Liénard system is a second-order ordinary differential equation that goes as
where [^8] Often it is sufficient to have $f(x),g(x)\in C^{2}(U,\mathbb{R})$, where $U\subset\mathbb{R}$. $f(x),g(x)\in C^{\infty}(\mathbb{R},\mathbb{R})$ can be suitably chosen. Interesting choices for $f(x)$ and $g(x)$ include $f(x)=1$ and $g(x)=x$, which is just the damped linear oscillator, while the choice $f(x)=(1-x^{2})$ and $g(x)=x$ gives the van der Pol oscillator VdP , known to admit limit-cycle behavior due to the particular choice of $f(x)$ str . Another choice is $f(x)=1$ and $g(x)=x^{3}$, for which we have the linearly-damped (nonlinear) Duffing oscillator (see for example, Refs. mic ; Demina21 ). It is noteworthy that in any case with $f(x)\neq 0$, the system exhibits non-conservative dynamics because (9) does not stay invariant under the transformation $t\rightarrow-t$, namely, time reversal. Further, oscillatory dynamics can be obtained if $f(x)$ is an even function and if $g(x)$ is odd; it follows from the fact that the overall force (the second and third terms of (9)) should be odd under $x\rightarrow-x$ in order to support oscillations mic .
@@ III.1 Cheillini condition and nonstandard Lagrangians
Given a second-order differential equation, the inverse problem of finding the Lagrangian has been the subject of much investigation whittaker ; yan78 ; Leach1 ; Leach2 ; JLMLienard ; mitra24 (see also, Ref. carinenaintegrability ). In particular, for Liénard systems satisfying a certain integrability condition, one can find nonstandard Lagrangians from which they emerge as the Euler-Lagrange equation JLMLienard (see also, Refs. agp ; BGG ). The idea is to make use of the so-called Jacobi last multiplier which may be defined as in whittaker (see Appendix (A)).
In this manner, starting from an ordinary differential equation
@@
one defines the last multiplier $M$ as that which satisfies
@@
As has been discussed in Whittaker’s classic textbook whittaker , if a second-order differential equation such as (10) follows from the Euler-Lagrange equations, then the Lagrangian is related to the last multiplier as
$M=\frac{\partial^{2}L(x,\dot{x})}{\partial\dot{x}^{2}}.$ (12)
This allows one to determine the Lagrangian function for a given second-order differential equation, provided it admits a Lagrangian formalism.
For the Liénard system, a formal solution for the last multiplier is found to be
@@
We may define a new nonlocal variable $u$ as JLMLienard
$u=\dot{x}-\frac{g(x)}{\ell f(x)},$ (14)
with $\ell\neq-1$ which is determined from
@@
Then using Eqs. (14) and (15) we have $\dot{u}=\ell uf(x)$. In other words, if the condition (15) is true the Liénard system (9) can be expressed as the following system of first-order equations:
$\dot{u}=\ell uf(x),\quad\quad\dot{x}=u+W(x),$ (16)
where $W(x)=\ell^{-1}g(x)/f(x)$. Since we have $\dot{u}=\ell uf(x)$, using (13) we can write $M=u^{1/\ell}$, which, upon using (12) gives us (up to a gauge function which can be ignored here) $L\sim u^{\frac{2\ell+1}{\ell}}$. In terms of $x$ and $\dot{x}$ the Lagrangian reads as agp
$L(x,\dot{x})=\frac{\ell^{2}}{(\ell+1)(2\ell+1)}\left(\dot{x}-\frac{1}{\ell}\frac{g(x)}{f(x)}\right)^{{\frac{2\ell+1}{\ell}}},$ (17)
which is of the nonstandard kind [^9] It may be noted that one cannot at this stage set $f(x)=0$ in (14), (15), (16), or (17) to recover the conservative case. However, one can set $f(x)=0$ in (13) which gives $M=1$ and consequently, from (12) one has $L(x,\dot{x})=\frac{\dot{x}^{2}}{2}+J(x)\dot{x}+K(x)$, where $K(x)$ is the (conservative) scalar potential while $J(x)$ may be interpreted as a vector potential..
It is noteworthy that (15) represents what is known as the Cheillini condition, allowing one to recast the Liénard system in the form (16). Specifically, if $f(x)=ax^{\alpha}$, then (15) dictates that $g(x)$ must satisfy the following differential equation (see also, Refs. CG19 ; CG24 ):
@@
where $k\in\mathbb{R}$ is an integration constant. This gives the functional form of $g(x)$ so as to satisfy the Cheillini condition. In the cases where the Cheillini condition is satisfied the Lagrangian is given by (17).